OpenStudy (anonymous):

Find dy/dx for the equation x^3-2x^2y+4xy^2=20 its instructed to use the the product rule for the middle terms which is where things become fuzzy I have 3x^2(product rule)=0 I really could use some help

OpenStudy (anonymous):

Considering \[\frac{dy}{dx}=y'\]\[x^3-2x^2y+4xy^2=20\]Let's make the derivative for each member Exponential rule\[x^3=3x^2\]Product rule\[-2x^2y=-2*(x^2y'+2yx)\]Product rule again\[4xy^2=4*(2xyy'+y^2)\]Derivativo of constant\[20=0\]Pull everything together\[3x^2-2x^2y'-4yx+8xyy'+4y^2=0\]\[y'*(8xy-2x^2)=4xy-3x^2+4y^2\]The answer is:\[y'=\frac{4xy-3x^2+4y^2}{(8xy-2x^2)}\]

OpenStudy (anonymous):

thank you!